Suppose that a competitive firm’s marginal cost of producing output q is given by MC (q) = 3+2q. Assume that the market price for the firm’s product is birr 9.
What level of output should the firm produce to maximize profit?
Suppose the firm’s fixed cost is known to be birr 3, will the firm be earning positive, negative or zero profit? What is that profit?.
1
Expert's answer
2021-09-13T12:42:43-0400

Maximizing profits call for companies to set their marginal cost as marginal revenue. Based on the idea that this company operates in competitive environment, marginal revenue is similar to market price it encounters.


Therefore, the firm needs to set market price as the marginal cost in order to maximize profits.

"q = 3, or 9 = 3 +2q"

Producer surplus tends to be as the section below market price, which is $9.00 while above marginal cost curve, which is 3 + 2q. Since MC tends to be linear, the producer surplus tends to be triangular having base equal to 9-3 = $6. Triangle's height tends to be 3 such that P = MC.


Hence;


Producer surplus =(0.5)(6)(3) = 9


Producer surplus = $9.